1. What is the management fee (1) on a per available room basis and (2) as a percentage of total revenue for a 255-room hotel located in California that had an occupancy level of 62%, ADR of $84.53, a room revenue to total revenue % of 56.4%, and a gross operating profit % of 24.8%? The management fee agreement stipulated that the company would receive 3% of gross revenue, and 10% of gross operating profit.

               Please calculate annual room revenue (round to two decimal places) $ ___

2. Please use the information from Question 10 to calculate the Annual total revenue.

              Annual total revenue (round to two decimal places) $ ___

3. Please use the information from Question 10 to calculate the GOP.

               GOP (round to two decimal places) $ ____

4.Please use the information from Question 10 to calculate the Mgmt fee base fee.

               Mgmt fee base fee (round to two decimal places) $ ___

5. Please use the information from Question 10 to calculate the Mgmt fee incentive fee.

           Mgmt fee incentive fee (round to two decimal places) $ ___
Answer #1

Answer 1:

Given that:

Number of rooms =  255

Occupancy level = 62%,

Average daily rate (ADR) = $84.53

Annual room revenue = Number of rooms * Occupancy level * ADR * 365

= 255 * 62% * 84.53 * 365

= 4877930.445

= $4877930.45

Annual room revenue = $4,877,930.45

Answer 2:

Room revenue to total revenue % of 56.4%.

Annual room revenue = $4,877,930.445

Annual total revenue = 4877930.445 / 56.4% = $8,648,812.85

Annual total revenue = $8,648,812.85

Answer 3:

Gross operating profit in % = 24.8%

Gross operating profit = 8648812.845 * 24.8% = $2,144,905.59

Gross operating profit = $2,144,905.59

Answer 4:

The management fee agreement stipulated that the company would receive 3% of gross revenue, and 10% of gross operating profit.

Mgmt fee base fee = $8,648,812.85 * 3% = $259,464.39

Mgmt fee base fee = $259,464.39

Answer 5:

Mgmt fee incentive fee = $2,144,905.59 * 10% = $214490.56

Mgmt fee incentive fee = $214,490.56